OpenStudy (anonymous):

Ma-li can paint a certain room in about 5 hours, and alicia can paint the same room in about 4 hours. Approximately how many hours would it take Ma-li and alicia to paint that rooom if they worked together?

hero (hero):

\[\frac{M \times A}{M + A}\]

OpenStudy (anonymous):

can you explain that a little more please?

hero (hero):

That's the working together formula. Just input the values M = Mali's hours to complete A = Alicia's hours to complete

OpenStudy (anonymous):

so it would be 5x4/5+4 =20/9=2.22

hero (hero):

Which is approximately 2 hours, right?

OpenStudy (anonymous):

yes :) this helped a lot, thank you!